OpenStudy (anonymous):

In a candy box, there are 2 times as many candies with nuts as candies without nuts. There are 12 candies in the box. How many candies with nuts are in the box? A 4 B 9 C 6 D 8

OpenStudy (anonymous):

I think so too but my dad is saying its A or D

OpenStudy (butterflydreamer):

in a candy box there are two times as many with nuts than without nuts. So let's say x + n (nuts) = 12 (since 12 is the total amount of candies) Then we know there are TWO times as many with nuts than without nuts. So then, n = 2x Substitute it into x + n = 12 x + 2x = 12 3x = 12 x = 4 Then we want to know how many has nuts so 12 - 4 = ?

OpenStudy (anonymous):

So Its 8 :D

OpenStudy (butterflydreamer):

yyep ^_^

OpenStudy (anonymous):

I told him and hes laughing saying how he was right lol Thank you :)

OpenStudy (butterflydreamer):

xD LOOOL. No problem :))!